OpenStudy (anonymous):

alonso runs a company called speedy house painters his workers can paint 3/4 f a house in one day. how many days would it take to paint 6 houses?

OpenStudy (anonymous):

3/4 foot(f) a house in one day?

OpenStudy (anonymous):

wht

OpenStudy (anonymous):

in the question what does that f means?

OpenStudy (anonymous):

f means of

OpenStudy (anonymous):

Let 'x' be the total part of the house. Total house = 6x. In one day they can paint 3/4 part of house.i.e.\[1day = \frac{ 3 }{ 4 }x\]\[\frac{ 4 }{ 3 }day = x\]So they can complete a house in 4/3 day. So they can complete 6 houses in \[\frac{ 4 }{ 3 } \times 6 = 8\]So it will take 8 days to complete 6 houses.

OpenStudy (anonymous):

thnks

OpenStudy (anonymous):

you are welcomed.